# NavRL: Learning Safe Flight in Dynamic Environments

### v1

@@ Abstract
Safe flight in dynamic environments requires autonomous unmanned aerial vehicles (UAVs) to make effective decisions when navigating cluttered spaces with moving obstacles. Traditional approaches often decompose decision-making into hierarchical modules for prediction and planning. Although these handcrafted systems can perform well in specific settings, they might fail if environmental conditions change and often require careful parameter tuning. Additionally, their solutions could be suboptimal due to the use of inaccurate mathematical model assumptions and simplifications aimed at achieving computational efficiency. To overcome these limitations, this paper introduces the NavRL framework, a deep reinforcement learning-based navigation method built on the Proximal Policy Optimization (PPO) algorithm. NavRL utilizes our carefully designed state and action representations, allowing the learned policy to make safe decisions in the presence of both static and dynamic obstacles, with zero-shot transfer from simulation to real-world flight. Furthermore, the proposed method adopts a simple but effective safety shield for the trained policy, inspired by the concept of velocity obstacles, to mitigate potential failures associated with the black-box nature of neural networks. To accelerate the convergence, we implement the training pipeline using NVIDIA Isaac Sim, enabling parallel training with thousands of quadcopters. Simulation and physical experiments [^1] Experiment video link: https://youtu.be/fhRxS--Rhkc show that our method ensures safe navigation in dynamic environments and results in the fewest collisions compared to benchmarks in scenarios with dynamic obstacles.
@@ I Introduction
Autonomous unmanned aerial vehicles (UAVs) are increasingly employed in a variety of applications, such as exploration [1], search and rescue [2], and inspection [3]. These applications often involve highly dynamic environments that require UAVs to actively avoid collisions. Previous methods typically rely on handcrafted algorithms and hierarchical modules for prediction and planning, resulting in overly complex systems with parameters that are difficult to tune. In contrast, reinforcement learning (RL) enables UAVs to automatically learn complex decision-making through extensive experience, offering greater adaptability across different settings and showing significant potential for safe collision avoidance with various obstacles. Therefore, exploring and developing RL-based navigation methods is crucial for enhancing UAV flight safety in dynamic environments.
@@
Developing an RL-based navigation method suitable for real-world deployment presents several challenges. First, since reinforcement learning involves training the robot through collision experiences, the learning process must occur in simulated environments. This creates a sim-to-real transfer issue due to the gap between simulated and real-world sensory information, particularly with camera images. Previous works have attempted to address this issue by developing methods to reduce the gap between simulated and real-world environments [4][5][6] or by training the robot directly in real-world settings [7][8][9]. However, these approaches often require additional training steps and can be data-inefficient when trained in real-world scenarios. Second, even if the trained RL policy demonstrates satisfied performance, ensuring safety remains challenging due to the black-box nature of neural networks, necessitating an effective safety mechanism to prevent severe failures [10][11] [12][13]. Lastly, training a reinforcement learning policy requires a substantial amount of robot experience, and some previous methods [14][15][16] that collect data using a single robot often result in slow convergence speeds due to limited data diversity and reduced parallel exploration opportunities.
@@
  • NavRL Navigation framework: This work introduces a novel reinforcement learning-based UAV navigation system to ensure safe flight in dynamic environments.
@@ II Related Work
There has been extensive research on UAV navigation in dynamic environments, with most approaches relying on rule-based methods that use handcrafted algorithms and optimization techniques [19][20][21]. While these solutions have proven successful in certain environments, they often result in complex systems that require meticulous parameter tuning when environmental conditions change. In contrast, learning-based methods learn complex decision-making logic from training data, reducing system complexity and offering better adaptability to diverse environments. We categorize learning-based navigation methods primarily into supervised learning-based and reinforcement learning-based approaches.
@@
Supervised learning-based methods: Methods in this category train networks using labeled datasets for decision-making. Early approaches [22][8][23] involve deploying robots in real-world environments, collecting images, and manually labeling them with corresponding ground truth actions. Similarly, some methods [7][24][25] predict the safety of input images rather than directly outputting the final decision and then use handcrafted algorithms to control the robot. The methods mentioned above require manually labeled real-world data and often suffer from limited generalization ability due to insufficient data. Loquercio et al. [26] train the network on an autonomous driving dataset to achieve navigation, benefiting from its extensive data volume. Jung et al. [27] use a learning-based detector to enable a drone to pass through gates in racing, while in [28], iterative learning control is applied to reduce tracking error in this context. Recently, a series of works on foundation models for visual navigation has been developed using a vast amount of real-world experience from various robots [29][30][31]. Maalouf et al. [32] explore optimal neural network structures for lightweight quadcopters, and Simon et al. [33] demonstrate collision avoidance capabilities for drones equipped with monocular cameras using a learning-based depth estimation method [34]. A novel imperative learning approach based on semantic images is proposed in [35] to achieve semantically-aware local navigation.
Reinforcement learning-based methods: Compared to supervised learning-based methods, reinforcement learning-based approaches benefit from the abundant data in simulated environments. Some methods utilizing deep Q-learning [14][15][36][37] or value network learning [38][39] have demonstrated successful navigation; however, these methods are constrained to discrete action spaces, which can lead to suboptimal performance. In contrast, policy gradient-based methods [40][41][42][16], which use an actor-critic network structure, enable robot control in continuous action spaces and have become increasingly popular. Recent work using imitation learning [43] has demonstrated impressive agile flight in real-world environments. Moreover, Kaufmann et al. [44] leverage a policy trained in simulation to outperform human champions in drone racing, with several studies focusing on enhancing control performance [45][46]. Meanwhile, Song et al. [6] develop a vision-based network by distilling knowledge from a teacher policy with full privileged information, while Xing et al. [47] apply contrastive learning to improve the encoding of input images. To further ensure safety, a recovery policy using a reach-avoid network is introduced in [48] to prevent failures. Similarly, Kochdumper et al. [13] utilize reachability analysis to project unsafe policy actions to safe regions. However, their approach requires a precomputed reachability set, with computation costs that can grow exponentially with the action dimensions.
@@ III-A Obstacle Perception System
Static and dynamic obstacles are handled separately by our perception system due to their differing properties, as illustrated in Fig. 2. Static obstacles, which can have arbitrary shapes and sizes, are represented more accurately using a discretized format, such as an occupancy voxel map. Conversely, dynamic obstacles, typically modeled as rigid bodies, are represented using bounding boxes with estimated velocity information. For the perception of static obstacles, we create a 3D occupancy voxel map with a fixed memory size, determined by the maximum number of voxels suitable for the environment. The occupancy data for each voxel is then stored in a pre-allocated array. This design enables us to access occupancy information with constant time complexity ($\mathcal{O}(1)$). At each time step, we recursively update the log probability of occupancy for each voxel based on the latest depth image and clear the occupancy data for dynamic obstacles by iterating through their detected bounding boxes. It is important to note that this static occupancy map can be generated on the fly without the need for any prebuilt data.
@@
The dynamic obstacle velocity is estimated by the tracking module, which operates in two stages: data association and state estimation. In the data association stage, the goal is to establish correspondences between detected obstacles across consecutive time steps. To minimize detection mismatches, we construct a feature vector for each obstacle that includes its position, bounding box dimensions, point cloud size, and point cloud standard deviation, and then determine matches based on the similarity scores between potential matching obstacles. In the state estimation stage, a Kalman filter is used to estimate obstacle velocities, with a constant acceleration model applied to account for changes in velocity over time. It is worth noting that while the perception system in this framework is based on a camera, other sensors can also be used if they provide a similar obstacle representation.
@@ III-B Reinforcement Learning Formulation
The objective of our framework is to generate control actions based on the robot’s available information and continuously apply them to ensure safe navigation and collision avoidance over time. Considering the challenges of sim-to-real transfer and the high sample requirements in reinforcement learning, it is essential to design suitable state, action, and reward representations that effectively address these issues while achieving safe navigation. This section outlines the detailed formulation of these RL components.
@@
Compared to dynamic obstacles, static obstacles are represented as map voxels, which cannot be directly input into neural networks. Therefore, we perform 3D ray casting from the robot’s position against the map. In Fig. 3a, rays are cast horizontally in all directions within a maximum range using a user-defined ray casting angle interval. Similarly, Fig. 3b illustrates the same operation in the vertical plane. For each vertical ray casting angle $\theta_{i}$, the lengths of all rays in the horizontal plane are recorded into a vector $R_{\theta_{i}}$. Any ray that exceeds the maximum range is assigned the maximum range length. The representation of static obstacles is constructed by stacking the ray length vectors for all vertical ray angles:
@@
Figure 3: Illustration of map ray casting. Only rays within the maximum range are shown. (a) A top-down view of horizontally cast rays with a 360-degree casting angle. (b) A 3D side view displaying the vertically cast rays.
@@
b) static safety reward $r_{ss}$: The static safety reward ensures that the robot keeps a safe distance from static obstacles. Given the static obstacle states in Eqn. 4, it is defined as:
@@ III-C Network Design and Policy Training
Given that our state representation consists of multiple components, a preprocessing step is necessary before inputting the data into the RL policy network. Both static and dynamic obstacles are represented as 2D matrices, so we utilize convolutional neural networks (CNN) to extract their features and transform them into 1D feature embeddings, as illustrated in Fig. 2. These embeddings are then concatenated with the robot’s internal states to form the complete input feature for the policy and value networks. We employ the Proximal Policy Optimization (PPO) algorithm [17] to train the actor (policy) and critic (value) networks, both of which are implemented using Multi-Layer Perceptrons. The training process is conducted in NVIDIA Isaac Sim, where we implement parallel training by collecting data from thousands of quadcopters simultaneously. The training environment features a forest-like setting with both static and dynamic obstacles. Each robot is spawned at a random location with a random goal and is reset either upon colliding with an obstacle or at the end of the episode. To improve learning efficiency, we adopt a curriculum learning strategy. The environment initially starts with a relatively low obstacle density, which is gradually increased as the success rate surpasses a specified threshold. Our experiments show that this approach allows the RL policy to achieve a higher navigation success rate in complicated environments.
@@ III-D Policy Action Safety Shield
Due to the black-box nature of the neural network, we designed a safety shield mechanism to prevent severe failures caused by the trained policy. Given that our policy outputs velocity commands for robot control, we incorporate the concept of velocity obstacles (VO) [18] to evaluate the safety of the policy’s actions and correct them if they could result in collisions. The velocity obstacle represents the set of robot velocities that would result in a collision within a specified future time horizon. In the scenario depicted in Fig. 5a, where the robot encounters obstacles, the corresponding velocity obstacle regions are visualized in Fig. 5b. In this example, there are two obstacles moving at different velocities, each associated with a velocity obstacle region based on their relative positions and velocities with respect to the robot. Each velocity obstacle region is a combination of a circular area centered at points C1 or C2, with a radius proportional to the combined sizes of the robot and the obstacle, and a cone area extending from the origin, excluding the dotted lines. The relative velocities of the robot to Obstacle 1 and Obstacle 2 (shown as red and green arrows) lie within the velocity obstacle regions, indicating future collisions. Inspired by [52], we adopt a similar approach to compute the minimum changes in velocity ($\Delta V_{1}$ and $\Delta V_{2}$) required to exit velocity obstacle regions. Given the current RL policy action $V_{rl}$ from the network, if this action is not within any of the velocity obstacle regions, we set the safe action $V_{safe}$ equal to the policy action. Otherwise, we formulate the following optimization to project the policy action into the safe region:
@@ IV Result and Discussion
To evaluate the proposed framework, we present our training results under different configurations and conduct simulation and physical flight tests in various environments. The policy was trained in NVIDIA Isaac Sim on a NVIDIA GeForce RTX 4090 GPU for around 10 hours. The maximum velocity of the robot is set to 2.0 m/s. The system is implemented in ROS, with the RL policy in Python and the policy action safety shield in C++. The simulation experiments are conducted on the RTX 4090 desktop, while computations for the physical flights are performed on our quadcopter’s onboard computer (NVIDIA Jetson Orin NX). An Intel RealSense D435i camera is utilized for static and dynamic obstacle perception, and the LiDAR Inertial Odometry (LIO) algorithm [53] is adopted for accurate robot state estimation.

### v2

@@ Abstract
Safe flight in dynamic environments requires unmanned aerial vehicles (UAVs) to make effective decisions when navigating cluttered spaces with moving obstacles. Traditional approaches often decompose decision-making into hierarchical modules for prediction and planning. Although these handcrafted systems can perform well in specific settings, they might fail if environmental conditions change and often require careful parameter tuning. Additionally, their solutions could be suboptimal due to the use of inaccurate mathematical model assumptions and simplifications aimed at achieving computational efficiency. To overcome these limitations, this paper introduces the NavRL framework, a deep reinforcement learning-based navigation method built on the Proximal Policy Optimization (PPO) algorithm. NavRL utilizes our carefully designed state and action representations, allowing the learned policy to make safe decisions in the presence of both static and dynamic obstacles, with zero-shot transfer from simulation to real-world flight. Furthermore, the proposed method adopts a simple but effective safety shield for the trained policy, inspired by the concept of velocity obstacles, to mitigate potential failures associated with the black-box nature of neural networks. To accelerate the convergence, we implement the training pipeline using NVIDIA Isaac Sim, enabling parallel training with thousands of quadcopters. Simulation and physical experiments [^1] Experiment video link: https://youtu.be/EbeJW8-YlvI show that our method ensures safe navigation in dynamic environments and results in the fewest collisions compared to benchmarks.
Index Terms:
Aerial Systems: Perception and Autonomy, Reinforcement Learning, Collision Avoidance
@@ I Introduction
Autonomous unmanned aerial vehicles (UAVs) are widely used in applications like exploration [1], search and rescue [2], and inspection [3]. These tasks often occur in dynamic environments that require effective collision avoidance. Traditional methods rely on handcrafted algorithms and hierarchical modules, leading to overly complex systems with hard-to-tune parameters. In contrast, reinforcement learning (RL) allows UAVs to learn decision-making through experience, offering better adaptability and improved performance. Developing RL-based navigation methods is essential for enhancing UAV safety in dynamic environments.
@@
Developing an RL-based navigation method suitable for real-world deployment presents several challenges. First, since reinforcement learning involves training the robot through collision experiences, the learning process must occur in simulated environments. This creates a sim-to-real transfer issue due to the gap between simulated and real-world sensory information, particularly with camera images. Previous works have attempted to address this issue by developing methods to reduce the gap between simulated and real-world environments [4][5][6] or by training the robot directly in real-world settings [7][8][9]. However, these approaches often require additional training steps and can be data-inefficient when trained in real-world scenarios. Second, even if the trained RL policy demonstrates satisfactory performance, ensuring safety remains challenging due to the black-box nature of neural networks, necessitating an effective safety mechanism to prevent severe failures [10][11] [12][13]. Lastly, training a reinforcement learning policy requires a substantial amount of robot experience, and some previous methods [14][15][16] that collect data using a single robot often result in slow convergence speeds due to limited data diversity and reduced parallel exploration opportunities.
@@
  • The NavRL Navigation Framework: This work introduces a novel reinforcement learning-based UAV navigation system to ensure safe autonomous flight in dynamic environments. The NavRL navigation framework is made available as an open-source package on GitHub  [^2] Software available at: https://github.com/Zhefan-Xu/NavRL.
@@ II Related Work
Research on UAV navigation in dynamic environments often relies on rule-based methods with handcrafted algorithms [19][20][21], which can be complex and require careful tuning as conditions change. In contrast, learning-based methods reduce complexity and adapt better to varying environments. This section mainly categorizes learning-based navigation into supervised and reinforcement learning-based methods, while acknowledging the existence of other approaches.
Supervised learning-based methods: Methods in this category train networks using labeled datasets. Early approaches [8][22] deploy robots in real-world environments, collect images, and manually label them with ground truth actions. Similarly, some methods [7][23] predict the safety of input images rather than outputting the decision and then use handcrafted algorithms to control the robot.
@@
The methods mentioned above require manually labeled real-world data and often suffer from limited generalization ability due to insufficient data. Loquercio et al. [24] train the network on an autonomous driving dataset to achieve navigation, benefiting from its extensive data volume. Jung et al. [25] use a learning-based detector to enable a drone to pass through gates in racing, while in [26], iterative learning control is applied to reduce tracking error in this context. Works on foundation models for visual navigation has been developed using real-world experience from various robots [27][28][29]. Simon et al. [30] demonstrate collision avoidance capabilities for drones equipped with monocular cameras using a depth estimation method [31]. An imperative learning approach based on semantic images is proposed in [32] to achieve semantically-aware local navigation.
Reinforcement learning-based methods: Compared to supervised learning-based methods, reinforcement learning-based approaches benefit from the abundant data in simulation. Some methods utilizing Q-learning [14][15][33] or value learning [34][35] have demonstrated successful navigation. However, these methods are constrained to discrete action spaces, which can lead to suboptimal performance.
Policy gradient methods [16][36], using an actor-critic structure, have gained popularity for enabling robot control in continuous action spaces. Kaufmann et al. [37] trained a policy in simulation to outperform human champions in drone racing, with other studies enhancing control performance [38][39]. Song et al. [6] developed a vision-based network by distilling knowledge from a teacher policy with privileged information, while Xing et al. [40] applied contrastive learning to improve image encoding. In [41], a comprehensive reinforcement training pipeline is presented for UAV-related tasks.
To ensure safety, a recovery policy using a reach-avoid network is introduced in [42] to prevent failures. Similarly, Kochdumper et al. [13] utilize reachability analysis to project unsafe policy actions to safe regions. However, their approach requires a precomputed reachability set, with computation costs scaling exponentially with action dimensions.
Most existing methods are designed for navigation and collision avoidance in static environments. Some methods, such as [36], demonstrate dynamic obstacle avoidance but do not address collision avoidance in complex static scenarios. The method in [42] can handle both static and dynamic obstacles but may perform suboptimally by treating dynamic obstacles as static. Recent safety-focused approaches are either computationally expensive or require extensive training. These challenges, combined with the difficulty of sim-to-real transfer, motivate us to propose a framework that ensures safe navigation while avoiding both static and dynamic obstacles.
@@ III-A Obstacle Perception System
Static and dynamic obstacles are handled separately by our perception system due to their different properties, as illustrated in Fig. 2. Static obstacles, which can have arbitrary shapes and sizes, are represented more accurately using a discretized format, such as an occupancy voxel map. Conversely, dynamic obstacles, typically modeled as rigid bodies, are represented using bounding boxes with estimated velocity information. The rationale for using different representations for static and dynamic obstacles lies in their distinct characteristics. Static obstacles sometimes have irregular geometries, making a discrete representation more suitable for accurate shape description. In contrast, dynamic obstacles, typically rigid bodies, require a representation that captures their states as a whole entity. Additionally, updating the occupancy map with dynamic obstacles can introduce noise and latency, further justifying the need for separate representations. For the perception of static obstacles, we create a 3D occupancy voxel map with a fixed memory size, determined by the maximum number of voxels suitable for the environment. The occupancy data for each voxel is then stored in a pre-allocated array. This design enables us to access occupancy information with constant time complexity ($\mathcal{O}(1)$). At each time step, we recursively update the log probability of occupancy for each voxel based on the latest depth image and clear the occupancy data for dynamic obstacles by iterating through their detected bounding boxes. It is important to note that this static occupancy map can be generated on the fly without the need for any prebuilt data.
@@
The dynamic obstacle velocity is estimated by the tracking module, which operates in two stages: data association and state estimation. In the data association stage, the goal is to establish correspondences between detected obstacles across consecutive time steps. To minimize detection mismatches, we construct a feature vector for each obstacle that includes its position, bounding box dimensions, point cloud size, and point cloud standard deviation, and then determine matches based on the similarity scores between potential matching obstacles. In the state estimation stage, a Kalman filter is used to estimate obstacle velocities, with a constant acceleration model applied to account for changes in velocity over time. It is worth noting that while the perception system in this framework is based on a camera, other sensors can also be used if they provide a similar obstacle representation. The dynamic obstacle detection and tracking in the perception system is based on the approach presented in [43], where readers can find further details and implementation specifics.
@@ III-B Reinforcement Learning Formulation
The navigation task is formulated as a Markov Decision Process (MDP) defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the state space (robot’s internal and sensory data), $A$ is the action space, the transition function $P(s_{t+1}|s_{t},a_{t})$ models environment dynamics, and the reward function $R(s_{t},a_{t})$ encourages goal-reaching behaviors while penalizing collisions and inefficient actions. The goal is to learn an optimal policy $\pi^{*}(a_{t}|s_{t})$ that maximizes the expected cumulative reward:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t})\right],$ (1)
where $\gamma\in[0,1]$ is the discount factor for future rewards. This formulation enables the use of reinforcement learning to train the robot for safe navigation in dynamic environments.
@@
Compared to dynamic obstacles, static obstacles are represented as map voxels, which cannot be directly input into neural networks. Therefore, we perform 3D ray casting from the robot’s position against the map. In Fig. 3a, rays are cast horizontally in all directions within a maximum range using a user-defined ray casting angle interval. Similarly, Fig. 3b illustrates the same operation in the vertical plane. For each diagonal ray casting angle $\theta_{i}$ in the vertical plane, the lengths of all rays in the horizontal plane are recorded into a vector $R_{\theta_{i}}$. Any ray exceeding the maximum range is assigned a length equal to the maximum range plus a small offset, allowing obstacle absence to be identified. The representation of static obstacles is constructed by stacking the ray length vectors for all diagonal ray angles in the vertical planes:
@@
Figure 3: Illustration of map ray casting. Only rays within the maximum range are shown. (a) A top-down view of horizontally cast rays with a 360-degree casting angle. (b) A side view displaying rays in the vertical planes.
@@
b) static safety reward $r_{ss}$: The static safety reward ensures that the robot keeps a safe distance from static obstacles. Given the static obstacle states in Eqn. 5, it is defined as:
@@ III-C Network Design and Policy Training
Given that our state representation consists of multiple components, a preprocessing step is necessary before inputting the data into the RL policy network. Both static and dynamic obstacles are represented as 2D matrices, so we utilize convolutional neural networks (CNN) to extract their features and transform them into 1D feature embeddings, as illustrated in Fig. 2. These embeddings are then concatenated with the robot’s internal states to form the complete input feature for the policy and value networks. We employ the Proximal Policy Optimization (PPO) algorithm [17] to train the actor (policy) and critic (value) networks, both of which are implemented using multi-layer perceptron. The training process is conducted in NVIDIA Isaac Sim, where we implement parallel training by collecting data from thousands of quadcopters simultaneously. The training environment features a forest-like setting with both static and dynamic obstacles. Each robot is spawned at a random location with a random goal and is reset either upon colliding with an obstacle or at the end of the episode. To improve learning efficiency, we adopt a curriculum learning strategy. The environment initially starts with a relatively low obstacle density, which is gradually increased as the success rate surpasses a specified threshold. Our experiments show that this approach allows the RL policy to achieve a higher navigation success rate in complicated environments.
@@ III-D Policy Action Safety Shield
Due to the black-box nature of the neural network, we designed a safety shield mechanism to prevent severe failures caused by the trained policy. Given that our policy outputs velocity commands for robot control, we incorporate the concept of velocity obstacles (VO) [18] to evaluate the safety of the policy’s actions and correct them if they could result in collisions. The velocity obstacle represents the set of robot velocities that would result in a collision within a specified future time horizon. In the scenario depicted in Fig. 5a, where the robot encounters obstacles, the corresponding velocity obstacle regions are visualized in Fig. 5b. In this example, there are two obstacles moving at different velocities, each associated with a velocity obstacle region based on their relative positions and velocities with respect to the robot. Each velocity obstacle region combines a circular area centered at C1 or C2, with a radius equal to the sum of the robot and obstacle sizes plus user-defined safe space, and a cone area extending from the origin, excluding the dotted lines. The relative velocities of the robot to Obstacle 1 and Obstacle 2 (shown as red and green arrows) lie within the velocity obstacle regions, indicating future collisions. Inspired by [46], we adopt a similar approach to compute the minimum changes in velocity ($\Delta V_{1}$ and $\Delta V_{2}$) required to exit velocity obstacle regions. The policy action produced by the network is defined as the velocity $V_{rl}$. If this action is not within any of the velocity obstacle regions, we set the safe action $V_{safe}$ equal to the policy action. Otherwise, we formulate the following optimization to project the policy action into the safe region:
@@ IV Result and Discussion
To evaluate the proposed framework, we present our training results under different configurations and conduct simulation and physical flight tests in various environments. The policy was trained in NVIDIA Isaac Sim on a NVIDIA GeForce RTX 4090 GPU for around 10 hours. The maximum velocity of the robot is set to 2.0 m/s. The simulation experiments are conducted on the RTX 4090 desktop, while computations for the physical flights are performed on our quadcopter’s onboard computer (NVIDIA Jetson Orin NX). An Intel RealSense D435i camera is utilized for static and dynamic obstacle perception, and the LiDAR Inertial Odometry (LIO) algorithm [47] is adopted for accurate robot state estimation. The static and dynamic feature extractors use 3-layer convolutional neural networks, outputting embeddings of sizes 128 and 64, respectively. The policy network consists of a two-layer multi-layer perceptron with a PPO clip ratio of 0.1. The ADAM optimizer is used with a learning rate of $5\times 10^{-4}$. The reward discounting factor is set to 0.99.